OpenStudy (anonymous):

how do i write y=x2-8x-1 in vertex form

OpenStudy (anonymous):

vertex The point on an angle where the two sides intersect.

OpenStudy (anonymous):

The vertex form of a quadratic function is given by f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola.

OpenStudy (bluebeta):

y=(x−4)2−17

OpenStudy (bluebeta):

the 2 should be ^2

OpenStudy (bluebeta):

y=(x-4)^2-17 or y+17=(x-4)^2

OpenStudy (tkhunny):

Have you considered Completing the Square? Think on this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (bluebeta):

who reported me?!